Grade 12Algebra

Solve 6x^{6}-25x^{5}+31x^{4}-31x^{2}+25x-6=0 of theory of equations

Y. Vijay
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear student
 
 
6x^6 - 25x^5 + 31x^4 - 31x^2 +25x - 6 = 0 
6x^6-6 -25x^5+25x+31x^4-31x^2=0 
6(x^6-1)-25x(x^4-1)+31x^2(x^2-1)=0 
6(x^2-1)(x^4+x^2+1)-25x(x^2-1)(x^2+1)+... 
(x^2-1){6x^4+6x^2+6-25x^3-25x+31x^2}=0 
(x^2-1){6x^4-25x^3+37x^2-25x+6}=0 
i.e, either, x^2-1=0, 0r, (6x^4-25x^3+37x^2-25x+6)=0 
x^2=1 gives roots x=1,-1 the other roots can be found out by the other part equating to zero, which is likely to return imaginary roots
 
 
 
 
Regards
Arun (askIITians forum expert)